# Opinion Clustering under the Friedkin-Johnsen Model: Agreement in Disagreement

### v1

@@ 5 Topology-based opinion clustering
Lemma 3 establishes that an LTP agent and the agents it persuades form a disjoint group in the network. However, there can be certain agents who are neither an LTP agent nor are persuaded by any LTP agent. Excluding them, we can uniquely partition the remaining agents into disjoint groups, with each group consisting of one LTP agent and its persuaded agents. The following results presents the relation between the final opinions of agents within each such group.
Lemma 5.2.
Consider a weakly connected network $\mathcal{G}$ with matrix $R$ derived using eqn. (3). Let $p$ be an LTP agent and $q\in\mathcal{N}_{p}$. If $\alpha$ is the union of set $\{p,q,n+1,....,n+m\}$ and all the oblivious agents in iSCC in $\mathcal{G}$ and $\alpha^{c}=[n+m]\setminus\alpha$ such that $\alpha^{c}\neq\phi$. Then,
Proof 5.3.
By definition of $\alpha$, we get $\alpha^{c}\subset[n]$. From the construction of $R$, it follows that $R[\alpha^{c}]=I-(I-\beta[\alpha^{c}])W[\alpha^{c}]$. If $\rho((I-\beta[\alpha^{c}])W[\alpha^{c}])<1$, then identity (5) follows by Neumann series. Let $M=(I-\beta[\alpha^{c}])W[\alpha^{c}]$. To prove that $\rho(M)<1$, it suffices to show that (I) $M$ is row substochastic, (II) for each $i^{th}$ row in $M$ with row-sum is $1$ there exists a $j^{th}$ row with row-sum less than $1$ such that there is a path from $j$ to $i$ in the associated digraph $\mathcal{G}(M)$. Then, by Theorem 6.37 in [18], $\rho(M)<1$.
Step-I: $M$ can be row-substochastic because: (i) a node $i\in\alpha^{c}$ whose in-neighbour is $p,q$ or an oblivious agent in iSCC satisfies $\sum_{j\in\alpha^{c}}w_{ij}<1$, (ii) each stubborn agent $i$ has the corresponding row-sum at most equal to $(1-\beta_{i})<1$. In an arbitrary digraph, the following scenarios may occur:
Thus, there always exists a node in $\alpha^{c}$ that satisfies either (i) or (ii). It implies that $M$ is row substochastic.
Step-II: As discussed in Step-I, each oblivious agent has a path from an oblivious agent in iSCC in $\mathcal{G}$. It follows from (i) that this path traverses an oblivious agent in $\alpha^{c}$ whose row-sum in $M$ is less than $1$. A special case could occur if each path in $\mathcal{G}$ from the oblivious agents in iSCC to an oblivious agent $\tilde{j}\in\alpha^{c}$ either traverses $p$ or $q$. In this case, the in-neighbours of $p$ or $q$ in $\alpha^{c}$ who have row-sum in $M$ less than $1$ lie on this path. Hence, every oblivious agent in $\alpha^{c}$ with row sum $1$ has a path in $\mathcal{G}(M)$ from a node with row sum <1<1. A similar argument applies to the non-oblivious agents as well. Thus, statement II also holds and $\rho(M)<1$.
@@
Consider a network $\mathcal{G}$ of $n$ agents with $m$ stubborn agents and no oblivious agents. The opinions of agents in $\mathcal{G}$ are governed by the FJ model (2). If $p$ is an LTP agent, then opinions of $p$ and nodes in set $\mathcal{N}_{p}$ converge to the same final opinion and form an opinion cluster.
Proof 5.4.
Let $q\in\mathcal{N}_{p}$. To prove that the final opinion $x_{p}^{*}=x_{q}^{*}$, we reduce the matrix $R$ in eqn. (3) using the Kron reduction and examine the relation between $x_{p}^{*}$ and $x_{q}^{*}$ in the reduced set of linear equations (4). Since $\mathcal{G}$ does not have any oblivious agents, each non-stubborn agent has a path from at least one stubborn agent. From the discussion in Remark 2, we know that the Kron Reduction $R/\alpha^{c}$ is well-defined for $\alpha=\{p,q,n+1,...,n+m\}\supset\{n+1,...,n+m\}$. Next, we evaluate the Schur complement $R/\alpha^{c}$ given by eqn. (1).
The following structural properties of the associated network $\mathcal{G}(R)$ will be use in determining $R/\alpha^{c}$,
  1. 1.
    Each walk from $s\in\mathcal{I}_{q}$ to $q$ in $\mathcal{G}$ traverses $p$ for all $s\in\mathcal{I}_{q}$. Thus, by construction of $\mathcal{G}(R)$, each walk that from $n+i$ to $q$ in $\mathcal{G}(R)$ will traverse $p$ for all $i\in[m]$.
  2. 2.
    The nodes n+1,,n+mn+1,...,n+m in $\mathcal{G}(R)$ are sources.
Property 1) implies that $r_{q,n+i}=0$ for all $i\in[m]$ because node $n+i$ cannot be an in-neighbour of $q$ in $\mathcal{G}(R)$. Further, property 2) implies that the rows of $R$ indexed $\{n+1,...,n+m\}$ have all entries as zero. Thus,
$R[\alpha]=\begin{bmatrix}r_{p,p}&r_{p,q}&r_{p,n+1}&...&r_{p,n+m}\\ r_{q,p}&r_{q,q}&0&...&0&\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$.
Let $F:[m+2]\to\alpha$ be a mapping such that $F(1)=p,F(2)=q$ and $F(j+2)=n+j$ for $j\in[m]$. This mapping relates the rows (and columns) indices of $Y$ (and $R/\alpha^{c}$) with the agents in set $\alpha$. Then, from eqn. (6), the entry $y_{ij}$ of $Y$ is non-zero only if there exists a walk in $\mathcal{G}(R)$ from $F(j)$ to $F(i)$ such that each node on this walk (except $F(i)$ and $F(j)$) belongs to $\alpha^{c}$. By property 1), $y_{2j}=0$ for all $j\in\{3,...,m+2\}$ because each walk from n+j2n+j-2 to $q$ in $\mathcal{G}(R)$ always passes through $p\in\alpha$. Consequently, $R/\alpha^{c}=R[\alpha]-Y$ has the following form:
$R/\alpha^{c}=\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&r_{1,3}^{1}&...&r_{1,m+2}^{1}\\ r_{2,1}^{1}&r_{2,2}^{1}&0&...&0\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$ where $r_{i,j}^{1}$ gives the entry of matrix $R/\alpha^{c}$ for $i,j\in[m+2]$. From Lemma 1, we know that $R/\alpha^{c}$ is also a Laplacian matrix and the row-sum of each of its rows equals 0. Since the second row of $R/\alpha^{c}$ has only two non-zero entries, thus, $r_{2,1}^{1}=-r_{2,2}^{1}$. Moreover, by eliminating the states $\alpha^{c}$ in the steady state eqns. (3), we get the following reduced system:
$\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&r_{1,3}^{1}&...&r_{1,m+2}^{1}\\ -r_{2,2}^{1}&r_{2,2}^{1}&0&...&0\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}\begin{bmatrix}x_{p}^{*}\\ x_{q}^{*}\\ \mathbf{x}_{s}(0)\end{bmatrix}=\mathbb{0}$ (7)
Simplifying eqn. (7) shows that $x_{p}^{*}=x_{q}^{*}$. Since $q$ is any node in $\mathcal{N}_{p}$, this holds for all nodes in $\mathcal{N}_{p}$.
Theorem 1 establishes that each LTP agent $p$ and the agents in $\mathcal{N}_{p}$ collectively form an opinion cluster in the final opinion. Importantly, the formation of an opinion cluster depends on the existence of an LTP agent in a network, independent of the edge weights and the stubbornness of the stubborn agents. The Example 1 illustrates the result in Theorem 1 as the LTP agent $3$ forms an opinion cluster with agents in $\mathcal{N}_{3}=\{4,5\}$ and LTP agent $6$ forms another opinion cluster with agent in $\mathcal{N}_{6}=\{1\}$. Additionally, it explains the formation of a new opinion cluster for the modified graph $\mathcal{\hat{G}}$. This occurs because the addition of edge $(1,5)$ results in a path from $6$ to $5$ that does not traverse $3$. Thus, $5\notin\mathcal{N}_{3}$ and forms a new opinion cluster in Fig. 1(d).
@@
An LTP agent can be a stubborn or a non-stubborn agent. However, it is crucial in ensuring that the nodes in its close vicinity in $\mathcal{N}_{p}$ form an opinion cluster. Thus, even a non-influential LTP agent impacts the final opinion formation, making it topologically influential.
The following result extends Theorem 1 to the networks containing the oblivious agents.
Theorem 2.
Consider a network $\mathcal{G}$ of $n$ agents with $m$ stubborn agents and at least one oblivious agent. The opinions of agents in $\mathcal{G}$ are governed by the FJ model (2). If each iSCC in $\mathcal{G}$ composed of the oblivious agents is aperiodic, then final opinions of an LTP agent $p$ and nodes in set $\mathcal{N}_{p}$ form an opinion cluster.
The proof is along the same lines as Theorem 1 and is omitted for brevity. By definition, if an LTP agent $p$ is oblivious, then the agents in $\mathcal{N}_{p}$ are also oblivious. Consequently, it follows from Theorem 2 that a set of non-influential oblivious agents can also form an opinion cluster. Since the opinion of an oblivious agent evolves according to DeGroot’s model and is not affected by stubborn behaviour, an LTP agent ensures the formation of opinion clusters even in DeGroot’s framework. Note that since nodes in $\mathcal{N}_{q}$ are always non-influential, the desired opinion clusters can be formed in DeGroot’s framework only in weakly connected networks.
@@
In [14], a network with $m$ stubborn agents and no oblivious agents has $m$ opinion clusters in the final opinion if the agents can be partitioned into subgraphs such that each subgraph contains only one stubborn agent and has a spanning tree with the stubborn agent as its root. Additionally, only the stubborn agent in a subgraph can receive information from the other subgraphs. An equivalent result exists for networks with oblivious agents as well. Note that it is assumed that each cycle in the network contains only one stubborn agent.
The topological conditions in [14] imply that all the paths to a node in a subgraph from the rest of the stubborn agents traverse the stubborn agent to the subgraph in which the node belongs. Thus, this stubborn agent is effectively the LTP agent for the agents within a subgraph that collectively form an opinion cluster. Hence, Theorems 1 and 2 extend the results presented in [14] to any weakly connected digraph.
Using the topological conditions in Theorems 1 and 2, a network can be designed to obtain a predefined number of opinion clusters with a desired subset of agents forming them. The following example highlights this property.
@@
Consider the network $\mathcal{G}$ with stubborn agents $8$ and $10$ in Fig. 4(a). The final opinion is desired to form opinion clusters with each cluster containing the agents within the dashed boxes in Fig. 4(a). To achieve desired clustering, we design $\mathcal{G}$ such that each dashed box (except the iSCC composed of oblivious agents that already form an opinion cluster) has an LTP agent and the remaining agents are persuaded by it. In Fig. 4(a), the nodes with green boundaries denote the LTP agents. For initial conditions from a uniform distribution over $[0,10]$ and stubbornness values in $[0,1)$, the opinions of agents in $\mathcal{G}$ under the FJ model form the desired opinion clusters as shown in Fig. 4(b)
@@ 6 Conclusions
The FJ model incorporates the individual biases as stubborn behaviour in the averaging-based opinion models. Due to varied biases, even the individuals in a closely connected group hold diverse opinions. Even then, a subgroup of individuals can still reach consensus, thereby, forming an opinion cluster. This work examines the role of the underlying network topology in the emergence of these clusters. To begin with, we define a topologically special agent called an LTP agent, and a set of non-influential agents that it persuades. Every path from an influential agent (a stubborn agent or an oblivious agent in an iSCC) to the persuaded agent always traverses the LTP agent. Using Kron-reduction, we establish in Theorems 1 and 2 that the final opinions of an LTP agent and the set of agents it persuades are equal. Thus, they form an opinion cluster in any arbitrarily-connected digraph under the FJ framework. Interestingly, such LTP agents lead to the formation of opinion clusters even under the DeGroot’s model (see Theorem 2).
Some additional key insights from our results are: (i) The presence of LTP agents always results in opinion clusters, independent of the edge-weights in the network and the stubbornness of the agents. (ii) Notably, an LTP agent may or may not be influential, but still shapes the final opinions of the agents it persuades. (iii) The notion of LTP agents generalises the topology-based conditions for opinion clustering in [14] to any arbitrary digraphs. (iv) By suitably placing the LTP agents and designing the network topology, we achieve any desired opinion clustering as demonstrated in Example 4.
@@
In future, we plan to examine the potential of LTP agents in desirably shaping the final opinions in real-world social networks.

### v2

@@ 4 Topology-based opinion clustering
Lemma 2 establishes that an LTP agent and the agents it persuades form a disjoint group in the network. Certain agents in $\mathcal{G}$ can exist who are neither an LTP agent nor are persuaded by any LTP agent. Excluding them, we can uniquely partition the remaining agents into disjoint groups, with each group consisting of one LTP agent and its persuaded agents.
4.1 The role of network topology on opinion formation
In this subsection, we present a framework that relates the final opinions of agents with the underlying network topology. Subsequently, we employ this framework to demonstrate that the topological relation between an LTP agent and the agent(s) it persuades leads to the formation of an opinion cluster.
Consider a network $\mathcal{G}$ of $n$ agents with opinions evolving under the FJ model (1). Let $\mathcal{G}$ have $m$ stubborn agents. Then, the final opinions satisfy: $\mathbf{x}^{*}=(I-\beta)W\mathbf{x}^{*}+\beta\mathbf{x}(0)$. We can rewrite these linear equations as follows:
where $\mathbf{x}_{s}=[x_{s_{1}}(0),...,x_{s_{m}}(0)]$ and $s_{1},s_{2},...,s_{m}\in[m]$ are the labels of the $m$ stubborn agents. The matrix $\eta=[\eta_{ij}]\in\mathbb{R}^{n\times m}$ has entry $\eta_{{i}j}=\beta_{i}$ if the stubborn agent $i$ is labelled as $s_{j}$, otherwise $\eta_{ij}=0$ for $i\in[n],j\in[m]$. Henceforth, we represent eqn. (2) as $Rz=\mathbb{0}$ where $R=[r_{i,j}]\in\mathbb{R}^{(n+m)\times(n+m)}$.
Remark 4.2.
The digraph $\mathcal{G}(R)$ can be induced from $R$ if an edge $(i,j)$ exists in $\mathcal{G}(R)$ when $r_{ji}<0$. The associated graph $\mathcal{G}(R)$ has $n+m$ nodes. Being consistent with the indexing of nodes in $\mathcal{G}$, the first $n$ nodes in $\mathcal{G}(R)$ are associated with $x_{i}^{*}$ for $i\in[n]$ and the nodes $n+1$ to $n+m$ are associated with the initial states of the $m$ stubborn agents. Each node $n+i$ forms a source in $\mathcal{G}(R)$ for $i\in[m]$. The source $n+i$ has only one outgoing edge to the node associated with the final opinion of the corresponding stubborn agent. For example, in Fig. 3, the initial opinion of stubborn agent $2$ is associated with node $n+1$ in $\mathcal{G}(R)$, thus, $n+1$ only has outgoing edge $(n+1,2)$.
Example 3
Consider the network $\mathcal{G}$ in Fig. 1(a) with stubborn agents $2$ and $6$. The corresponding network $\mathcal{G}(R)$, shown in Fig. 3, is derived from eqn. (2). Each node associated with the initial opinion of a stubborn agent is shown in turquoise.
Figure 3: The network $\mathcal{G}(R)$ derived from $\mathcal{G}$.
Remark 4.3.
Under Kron Reduction, we determine the Schur complement (4) of the loopy Laplacian matrix $Q$. Since $R$ is a Laplacian matrix, it can be reduced by Kron Reduction for a suitable choice of the subset of nodes $\alpha$. If $\mathcal{G}$ does not have any oblivious agents, then each node in $[n]$ has a path in $\mathcal{G}(R)$ from at least one source $\{n+1,...,n+m\}$. Thus, by Statement 1) in Lemma 5.1, the Schur complement $R/\alpha^{c}$ is well-defined for any $\alpha$ such that $\{n+1,...,n+m\}\subseteq\alpha\subseteq[n+m]$.
On the contrary, if $\mathcal{G}$ has an oblivious agent, then the sources $\{n+1,...,n+m\}$ do not have a path in $G(R)$ to this agent. For such networks, the set $\alpha$ must include the oblivious agents belonging to each iSCC and sources $\{n+1,...,n+m\}$. For this choice of $\alpha$, each $i\in\alpha^{c}$ has a path in $\mathcal{G}(R)$ from a node $j\in\alpha$. Thus, Statement 1) in Lemma 5.1 holds.
Consider a set of linear equations $M\mathbf{y}=\mathbb{0}$, where $\mathbf{y}\in\mathbb{R}^{p}$. Eliminating $y[\alpha^{c}]$ using Schur complement (4), yields the reduced system of equations: $M/\alpha^{c}\ \mathbf{y}[\alpha]=0$ (Pg. 18 in [21]). Similarly, we eliminate the states $z[\alpha^{c}]$ in the linear eqns. (2) using the Schur Complement to obtain,
where $\omega=\alpha\setminus\{n+1,...,n+m\}$. Thus, the Kron Reduction technique is useful in determining the relations between the final opinions of a subgroup of nodes in $\mathcal{G}$ and quantifying their dependence on the influential agents.
4.2 Main Results
Using the Kron Reduction technique, we present topology-based conditions under which the agents in a subgroup have equal final opinions and form an opinion cluster.
@@ 4.2 Main Results
Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents. Let Assumption 1 hold. The opinions of agents in $\mathcal{G}$ are governed by the FJ model (1). If $p$ is an LTP agent, then $p$ and the nodes in set $\mathcal{N}_{p}$ form an opinion cluster.
The proof is given in the Appendix.
In [15], the authors show that the final opinion of an agent depends on the paths (the weights of edges on the path) to it from the stubborn agents. However, it remains unclear what topological conditions on the paths from stubborn agents to distinct agents ensure that the final opinions of these agents converge to the same value. Theorem 1 addresses this gap by proving that the topological relation between an LTP agent and the agents it persuades leads to the formation of an opinion cluster. A major advantage of this result is that regardless of edge weights in the network $\mathcal{G}$ and the stubbornness values, an LTP agent $p$ and the agents in $\mathcal{N}_{p}$ collectively form an opinion cluster. Example 1 illustrates the result in Theorem 1 as the LTP agent $3$ and agents in $\mathcal{N}_{3}=\{4,5\}$ form an opinion cluster. Similarly, the LTP agent $6$ forms another opinion cluster with agent in $\mathcal{N}_{6}=\{1\}$.
@@
In [12], the authors derived the conditions to obtain $k\in\mathbb{N}$ opinion clusters in weakly connected digraphs, under the assumption that each cycle in the digraph has at most one stubborn agent. Theorem 1 generalises these conditions to form opinion clusters in any arbitrarily connected digraphs, thereby removing this restrictive assumption.
Moreover, in [12] an opinion cluster is formed by group of agents composed either (1) of a single stubborn agent with the rest being non-stubborn, or (2) of all the oblivious agents in the network $\mathcal{G}$ when $\mathcal{G}$ has a unique aperiodic iSCC of oblivious agents. Hence, a group containing only non-influential agents does not form any opinion cluster in [12]. In contrast, Theorem 1 establishes that an opinion cluster can be formed by any group of non-influential agents if one of them is an LTP agent $p$ and the rest are in $\mathcal{N}_{p}$ (e.g., the agents $3,4,5$ in Example 1).
@@
Consider an agent $p\in\mathcal{V}$. To verify if $p$ is an LTP agent, we first remove all the outgoing edges of $p$, which takes $O(n)$ time. Next, the modified $\mathcal{G}$ is traversed $m+t$ times using the Depth First Search (DFS), once from each of the $m$ stubborn agents and the $t$ iSCCs composed of oblivious agents. As one iteration of DFS has complexity $O(n+|\mathcal{E}|)$, this step requires $O((m+t)(n+|\mathcal{E}|))$ time, where $|\mathcal{E}|$ denotes the number of edges. If a node remains unvisited after all these traversals, then $p$ satisfies Condition (i) in Defn 2. Let $\mathcal{U}_{p}$ be the set of unvisited nodes. By repeating this step for all $p\in\mathcal{V}$, we identify all agents that satisfy Condition (i) and store them as $\mathcal{H}$. This step has a time complexity of $O((t+m)n(n+|\mathcal{E}|))$ which reduces to $O(n(n+|\mathcal{E}|))$ since $t+m\ll n$. If $\mathcal{H}\subseteq\mathcal{I}$, identifying all LTP agents requires $O(n(n+|\mathcal{E}|))$ time.
Next, we verify condition (ii) in Defn. 2 for each non-influential agent in $\mathcal{H}$. For an non-influential agent $p$, we check if there is a $c\in\mathcal{H}\setminus\{p\}$ such that $p\in\mathcal{U}_{c}$. This step requires $O(n^{2})$ time. Repeating it for all $p\in\mathcal{H}$ yields a worst-case complexity of $O(n^{3})$. Thus, the overall time complexity of identifying all LTP agents in $\mathcal{G}$ is $O(n^{3})$.
Using the topological conditions in Theorems 1, a network can be designed to obtain a predefined number of opinion clusters with each cluster formed by a desired set of agents.
Corollary 2.
Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents and $t\geq 0$ iSCCs with oblivious agents such that Assumption 1 holds. The opinions of agents in $\mathcal{G}$ are governed by the FJ model (1). Let $\mathcal{O}$ be the set of the oblivious agents that either belong to an iSCC or have paths from only one iSCC in $\mathcal{G}$. Then, $k\geq t+m$ opinion clusters form in $\mathbf{x}^{*}$ if the agents in set $\mathcal{V}\setminus\mathcal{O}$ can be partitioned into ktk-t disjoint groups with each group composed of an LTP agent and the nodes persuaded by it.
The proof of Corollary 2 is given in the Appendix. The following example illustrates the conditions presented in Corollary 2 to obtain the desired opinion clustering.
@@
Consider the network $\mathcal{G}$ with stubborn agents $8$ and $10$ in Fig. 4(a). The agents $1,2$ and $3$ are oblivious agents which together form an iSCC (highlighted by a dashed red circle); the remaining agents are non-oblivious. The non-oblivious agents in $\mathcal{G}$ form $4$ disjoint groups highlighted by the green dashed boxes in Fig. 4(a). Each group has an LTP agent (the node with green boundary) and the remaining agents in the group are persuaded by it. For initial conditions from a uniform distribution over $[0,10]$ and stubbornness values in $[0,1)$, the opinions of agents in $\mathcal{G}$ under the FJ model form $5$ opinion clusters as shown in Fig. 4(b). One cluster is formed by the oblivious agent belonging to the iSCC and the rest are formed by the agents belonging to the same group.
@@ 5 Conclusions
The FJ model incorporates the individual biases as stubborn behaviour in the averaging-based opinion models. Due to varied biases, even the individuals in a closely connected group hold diverse opinions. Even then, a subgroup of individuals can still reach consensus, thereby, forming an opinion cluster. This work examines the role of the underlying network topology in the emergence of these clusters. To begin with, we define a topologically special agent called an LTP agent, and a set of non-influential agents that it persuades. Every path from an influential agent (a stubborn agent or an oblivious agent in an iSCC) to the persuaded agent always traverses the LTP agent. Using Kron Reduction, we establish in Theorem 1 that an LTP agent and the set of agents it persuades form an opinion cluster in any arbitrarily-connected digraph under the FJ framework. Some additional key insights from our results are as follows: (i) The presence of LTP agents always results in opinion clusters, independent of the edge-weights in the network and the stubbornness of the agents. (ii) The notion of LTP agents generalises the topology-based conditions for opinion clustering in [12] to any arbitrary digraphs. (iii) By suitably placing the LTP agents and designing the network topology, we achieve any desired opinion clustering.
@@
In future, we plan to use the LTP agents for desirably shaping the final opinions in real-world social networks.
Appendix
Kron Reduction: Let $M\in\mathbb{R}^{p\times p}$ with index set $\alpha\subseteq[p]$ and $\alpha^{c}:=[p]\setminus\alpha$. If $M[\alpha^{c}]$ is non-singular, then the Schur complement of $M[\alpha^{c}]$ in $M$ is defined as:
Under Kron Reduction, the Schur complement of the loopy Laplacian matrix $Q$ is evaluated.
Lemma 5.1 ([18]).
  1. 1.
    $Q/\alpha^{c}$ is well defined ( i.e. $Q[\alpha^{c}]$ is invertible) if for each node $i\in\alpha^{c}$ there is a node $j\in\alpha$ such that a path $j\to i$ exists in $\mathcal{G}(Q)$. 111(Note that we adapt this condition to our framework because we consider $w_{ij}>0$ when $(j,i)\in\mathcal{E}$.)
  2. 2.
5.1 Proof of Theorem-1
Before moving to the proof of Theorem-1, the following Lemma needs to be stated.
Lemma 5.2.
Consider a weakly connected network $\mathcal{G}$ with $m\geq 1$ stubborn agents and matrix $R$ derived using eqn. (2). Let $p$ be an LTP agent and $q\in\mathcal{N}_{p}$. If $\alpha$ is the union of set $\{p,q,n+1,....,n+m\}$ and all the oblivious agents in iSCC in $\mathcal{G}$ and $\alpha^{c}=[n+m]\setminus\alpha$ such that $\alpha^{c}\neq\phi$. Then,
Proof 5.3.
By definition of $\alpha$, we get $\alpha^{c}\subset[n]$. From the construction of $R$, it follows that $R[\alpha^{c}]=I-(I-\beta[\alpha^{c}])W[\alpha^{c}]$. If $\rho((I-\beta[\alpha^{c}])W[\alpha^{c}])<1$, then identity (5) follows by Neumann series. Let $M=(I-\beta[\alpha^{c}])W[\alpha^{c}]$. To prove that $\rho(M)<1$, it suffices to show that (I) $M$ is row substochastic, (II) for each $i^{th}$ row in $M$ with row-sum is $1$ there exists a $j^{th}$ row with row-sum less than $1$ such that there is a path from $j$ to $i$ in the associated digraph $\mathcal{G}(M)$. Then, by Theorem 4.11 in [20])1, $\rho(M)<1$.
Step-I: $M$ is row-substochastic if: (i) a node $i\in\alpha^{c}$ whose in-neighbour is $p,q$ or an oblivious agent in iSCC satisfies $\sum_{j\in\alpha^{c}}w_{ij}<1$, (ii) node $i\in\alpha^{c}$ is a stubborn agent and the corresponding row-sum is at most $(1-\beta_{i})<1$. In an arbitrary digraph $\mathcal{G}$, the following scenarios may occur:
Thus, there always exists a node in $\alpha^{c}$ that satisfies either (i) or (ii). Hence, $M$ is row substochastic.
Step-II: As discussed in Step-I, each oblivious agent has a path from an oblivious agent in iSCC in $\mathcal{G}$. It follows from (i) that this path traverses an oblivious agent in $\alpha^{c}$ whose row-sum in $M$ is less than $1$. A special case could occur if each path in $\mathcal{G}$ from the oblivious agents in iSCC to an oblivious agent $\tilde{j}\in\alpha^{c}$ either traverses $p$ or $q$. In this case, the in-neighbour(s) of $p$ or $q$ in $\alpha^{c}$ who has row-sum in $M$ less than $1$ lie on this path. Hence, every oblivious agent in $\alpha^{c}$ with row sum $1$ has a path in $\mathcal{G}(M)$ from a node with row sum <1<1. A similar argument applies to the non-oblivious agents as well. Thus, statement II also holds and $\rho(M)<1$.
Proof of Theorem-1: Let $q\in\mathcal{N}_{p}$. To prove that the final opinion $x_{p}^{*}=x_{q}^{*}$, we reduce the matrix $R$ in eqn. (2) using the Kron Reduction and examine the relation between $x_{p}^{*}$ and $x_{q}^{*}$ in the reduced set of linear equations (3). Let the first $u$ agents in $\mathcal{G}$ be oblivious agents that form the iSCCs. The oblivious agents in each iSCC in $\mathcal{G}$ have consensus [9]. Hence, we consider $q$ is not an oblivious agent in an iSCC; otherwise, $p$ and $q$ belong to an iSCC that is known to achieve consensus.
Consider $\alpha$ to be $\{p,q,1,...,u,n+1,...,n+m\}$. From Remark 4.3, we know that the Kron Reduction $R/\alpha^{c}$ is well-defined. Next, we evaluate $R/\alpha^{c}$ given by eqn. (4). The following structural property of the associated network $\mathcal{G}(R)$ will be used in determining $R/\alpha^{c}$:
  1. P1:
    Each walk that exists from $s\in\mathcal{I}_{q}$ to $q$ in $\mathcal{G}$ traverses $p$ for all $s\in\mathcal{I}_{q}$. Thus, by construction of $\mathcal{G}(R)$, each walk that exists from $j$ to $q$ in $\mathcal{G}(R)$ will traverse $p$ for all $j\in\{[u],n+1,...,n+m\}$.
Property P1 implies that $r_{q,j}=0$ for all $j\in\{[u],n+1,...,n+m\}$ because node $j$ cannot be an in-neighbour of $q$ in $\mathcal{G}(R)$. Additionally, the nodes n+1,,n+mn+1,...,n+m in $\mathcal{G}(R)$ are sources. Thus, the corresponding rows of $R$ have all entries as zero. Therefore, $R[\alpha]$ takes the following form:
$R[\alpha]={\begin{bmatrix}r_{p,p}&r_{p,q}&R[p,[u]]&r_{p,n+1}&...&r_{p,n+m}\\ r_{q,p}&r_{q,q}&\mathbb{0}&0&...&0&\\ \mathbb{0}&\mathbb{0}&R[[u]]&\mathbb{0}&\cdots&\mathbb{0}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}}$.
Let $F:[{u}+m+2]\to\alpha$ be a mapping such that $F(1)=p,F(2)=q$, $F(j+2)=j$ for $j\in[u]$, and $F(u+2+j)=n+j$ for $j\in[m]$. This mapping relates the row (and column) indices of $Y$ (and $R/\alpha^{c}$) with the agents in set $\alpha$. Then, from eqn. (6), the entry $y_{ij}$ of $Y$ is non-zero only if there exists a walk in $\mathcal{G}(R)$ from $F(j)$ to $F(i)$ such that each node on this walk (except $F(i)$ and $F(j)$) belongs to $\alpha^{c}$. By property P1), $y_{2j}=0$ for all $j\in{\{3,...,u+m+2\}}$ because each walk from corresponding $F(j)$ to $q$ in $\mathcal{G}(R)$ always passes through $p\in\alpha$. Consequently, $R/\alpha^{c}=R[\alpha]-Y$ has the form:
$\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&R_{1}&r_{1,u+2}^{1}&...&r_{1,u+m+2}^{1}\\ r_{2,1}^{1}&r_{2,2}^{1}&\mathbb{0}_{u}&0&...&0\\ \mathbb{0}_{u\times 1}&\mathbb{0}_{u\times 1}&R[[u]]&\mathbb{0}_{u\times 1}&\cdots&\mathbb{0}_{u\times 1}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$}
where $r_{i,j}^{1}$ gives the entry of matrix $R/\alpha^{c}$ for $i,j\in[m+2]$. The matrix $R_{1}$ is the submatrix $R/\alpha^{c}[1,[u]]$ highlighting the corresponding non-zero entries. From Lemma 5.1, we know that $R/\alpha^{c}$ is also a Laplacian matrix and the row-sum of each of its rows equals 0. Since the second row of $R/\alpha^{c}$ has only two non-zero entries, thus, $r_{2,1}^{1}=-r_{2,2}^{1}$. Hence, by eliminating the states $\alpha^{c}$ in the steady state eqns. (2), we get $x_{p}^{*}=x_{q}^{*}$. Since $q$ is any node in $\mathcal{N}_{p}$, this holds for all nodes in $\mathcal{N}_{p}$. $\blacksquare$
5.2 Proof of Corollary 2
We classified the agents in $\mathcal{G}$ into sets $\mathcal{O}$ and $\mathcal{V}\setminus\mathcal{O}$. As $\mathcal{V}\setminus\mathcal{O}$ is union of tkt-k disjoint sets $\{i\}\cup\mathcal{N}_{i}$, where each $i\in\mathcal{L}$ is an LTP agent. Hence, ktk-t LTP agents are present in set $\mathcal{V}\setminus\mathcal{O}$. By Theorem 1, each LTP agent $i$ and the agents in $\mathcal{N}_{i}$ form an opinion cluster. Therefore, the agents in set $\mathcal{V}\setminus\mathcal{O}$ form ktk-t clusters. The set $\mathcal{V}\setminus\mathcal{O}$ contains the $m$ stubborn agents. Since stubborn agents persuaded by an LTP agent, each stubborn agent must be an LTP agent. Thus, $k-t\geq m$.
[…]
Let $C_{j}$ be the set of oblivious agents in $j^{th}$ iSCC and those oblivious agents only have a path from the $j^{th}$ iSCC. The opinions of agents in $C_{j}$ evolve as $\mathbf{y}_{j}(k)=W_{i}^{j}\mathbf{y}_{j}(k)$ with $\mathbf{y}_{j}$ is the vector that denotes the opinion of agents in this set. Since the associated digraph $G(W_{i}^{t})$ is a subgraph of $\mathcal{G}$ containing a unique aperiodic iSCC, the opinions of the agents converge to consensus (Theorem 5.1 in [20])1.
Hence, the $t$ subgroups, each composed of an iSCC of oblivious agents and those having a path only from the iSCC, form $t$ opinion clusters. Collectively, the oblivious agents in $\mathcal{O}$ and the agents in $\mathcal{V}\setminus\mathcal{O}$ form $k$ opinion clusters such that $k\geq t+m$.